# CMT: Mid-Training for Efficient Learning of Consistency, Mean Flow, and Flow Map Models

### v1

@@ 4.2 Mid-Training with CMT Improves Flow Map Post-Training
These results show that diffusion initialization is insufficient and underscore the importance of mid-training. The use of B/4 for pre-training and XL/2 for mid-training highlights a unique feature of CMT: its mid-training is architecture agnostic, as it directly learns a map aligned with the teacher sampler.
@@ 5 Theoretical Analysis
In mid–training, CMT learns a reliable proxy of the flow map, yielding a well–aligned initialization for post–training. We assess this effect by analyzing the CM flow map ${\bm{\Psi}}_{t\to 0}$ in the post–training stage; the same reasoning applies to the general flow map ${\bm{\Psi}}_{t\to s}$, such as MF.
Our goal is to theoretically quantify how closely the surrogate CM objective aligns with the oracle descent directions under different initialization schemes: CMT from mid-training (${\bm{\theta}}_{\mathrm{CMT}}$), a pre-trained diffusion model (${\bm{\theta}}_{\mathrm{DM}}$), and random initialization (${\bm{\theta}}_{\mathrm{rand}}$). More precisely, we consider the squared $\ell_{2}$ distance $d({\mathbf{x}},{\mathbf{y}})=\|{\mathbf{x}}-{\mathbf{y}}\|_{2}^{2}$ and a uniform weight $w(t)\equiv 1$. Let $\mathcal{L}_{\mathrm{oracle}\text{-}\mathrm{CM}}({\bm{\theta}})$ denote the oracle objective (Equation 2) and $\mathcal{L}_{\mathrm{CM}}({\bm{\theta}})$ the surrogate CM objective (Equation 3). We define the gradient bias as
$\mathcal{B}({\bm{\theta}}):=\big\|\nabla_{{\bm{\theta}}}\mathcal{L}_{\mathrm{oracle}\text{-}\mathrm{CM}}({\bm{\theta}})-\nabla_{{\bm{\theta}}}\mathcal{L}_{\mathrm{CM}}({\bm{\theta}})\big\|_{2}^{2}.$
We evaluate each scheme by $\mathcal{B}({\bm{\theta}})$ at its initial parameters, i.e., at ${\bm{\theta}}\in\{{\bm{\theta}}_{\mathrm{CMT}},{\bm{\theta}}_{\mathrm{DM}},{\bm{\theta}}_{\mathrm{rand}}\}$. A smaller $\mathcal{B}$ indicates that updates on $\mathcal{L}_{\mathrm{CM}}$ closely track those on $\mathcal{L}_{\mathrm{oracle}\text{-}\mathrm{CM}}$. To characterize robustness, we bound $\mathcal{B}$ in the worst case for each initialization scheme, which leads to the theorem below.
Theorem 5.1 (Informal Bias Comparisons).
Fix an error tolerance $\varepsilon>0$ and a small time step $\Delta t$. The deviation between the CM flow map gradient and the oracle gradient satisfies:
@@
    CMT : If $\mathcal{L}_{\mathrm{CMT}\text{-}\mathrm{CM}}({\bm{\theta}}_{\mathrm{CMT}})<\varepsilon$, for some ${\bm{\theta}}_{\mathrm{CMT}}$, then $\mathcal{B}({\bm{\theta}}_{\mathrm{CMT}})=\mathcal{O}(\varepsilon+\Delta t^{2})$.
@@
    Diffusion Model: If $\mathcal{L}_{\mathrm{DM}}({\bm{\theta}}_{\mathrm{DM}})<\varepsilon$, for some ${\bm{\theta}}_{\mathrm{DM}}$, then
    $\mathcal{B}({\bm{\theta}}_{\mathrm{DM}})=\mathcal{O}\left(\varepsilon+\Delta t^{2}+\mathbb{E}_{t}\big[\tfrac{\sigma_{t}^{2}}{\alpha_{t}^{2}}\big]\right)+\mathbb{E}_{t,{\mathbf{x}}_{t}}\left[\|{\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})-\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]\|_{2}^{2}\right].$
@@
    Random Initialization: For randomly initialized weights ${\bm{\theta}}_{\mathrm{rand}}$, $\mathcal{B}({\bm{\theta}}_{\mathrm{rand}})=\mathcal{O}(1)$.
From Theorem 5.1, we see that initializing the flow map model with CMT’s mid-training weights provides a strong starting point, close to the oracle. This is because the mid-training stage of CMT already offers a good proxy for the oracle flow map.
In contrast, initialization with a pre-trained diffusion model, as suggested by Geng et al. (2025b), incurs unavoidable sources of bias. First, it introduces a large constant term $\mathbb{E}_{t}\left[\tfrac{\sigma_{t}^{2}}{\alpha_{t}^{2}}\right]$. In EDM’s formulation ($\alpha_{t}=1$, $\sigma_{t}=t$), this becomes $\mathbb{E}_{t}\left[\tfrac{\sigma_{t}^{2}}{\alpha_{t}^{2}}\right]=\mathcal{O}(T^{2})$ if $t\sim\mathrm{Unif}[0,T]$. Second, it suffers from the inherent mismatch between the PF-ODE solution and the posterior mean, captured by $\mathbb{E}_{t,{\mathbf{x}}_{t}}\!\left[\|{\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})-\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]\|_{2}^{2}\right]$, which is generally nonzero. Together, these confirm our earlier claim that flow maps must learn large integrated jumps, whereas diffusion models encode only infinitesimal shifts, creating a fundamental mismatch that makes diffusion-based initialization fragile.
Random initialization is even less favorable: training can start arbitrarily far from the oracle, and the resulting lack of control typically leads to slow convergence. We present the complete and rigorous version of the theorem in Theorem F.1, which also covers the case of initializing with a trained consistency distillation model. This setting introduces an additional, uncontrollable discrepancy compared to CMT, and in practice, its training often requires extra ad-hoc techniques, further limiting its robustness relative to CMT. We refer to Theorem F.2 for a complementary discussion of gradient variance, and note that the bias terms are dominant (see Corollary F.3). Combining these ingredients, Theorem F.4 shows that CM-driven SGD with CMT initialization achieves the smallest excess risk and the lowest final error among all other initialization schemes.
@@ 6 Conclusion
Appendix A Related Work
Diffusion Model and ODE Samplers. Diffusion models and flow matching learn a time-conditioned vector field that transports a simple prior to the data distribution. In diffusion models, the associated PF-ODE shares the same marginals as the reverse SDE, which permits deterministic generation via ODE solving (Song et al., 2021; Lipman et al., 2023; Liu et al., 2023). Early accelerations such as DDIM instantiate a first-order (Euler-like) discretization of the PF-ODE, reducing sampling steps without retraining (Song et al., 2020). The EDM design space refines parameterization, noise schedules, and training targets to improve stability and enable aggressive few-step sampling (Karras et al., 2022). High-order samplers in the DPM-Solver family leverage a log-SNR reparameterization of time and exponential-integrator schemes (orders 2/3) to integrate the PF-ODE accurately with very few NFEs (Lu et al., 2022; 2025).
Few-Step Flow Map Models. Diffusion sampling is slow due to stepwise SDE/ODE integration. Early distillation methods (Salimans & Ho, 2022; Luhman & Luhman, 2021; Zheng et al., 2023) accelerate sampling by training a student to match a multi-step diffusion teacher with long jumps. CM (Song et al., 2023; Song & Dhariwal, 2024) instead learn a direct flow map via pairwise time consistency, later improved in training and extended to continuous time (Geng et al., 2025b; Lu & Song, 2025). Consistency trajectory models (CTM) (Kim et al., 2024; Lai et al., 2023) extend CM to learn flow maps between arbitrary points along the trajectory, but their training relies on adversarial objectives. Follow-up works (Frans et al., 2025; Boffi et al., 2024; Sabour et al., 2025) propose alternative parameterizations and losses to avoid adversarial components while improving few-step fidelity. Mean Flow (Geng et al., 2025a) attempts to train from scratch without adversarial loss, at the cost of heavy computation from Jacobian–vector product evaluations.
Relationship of Flow Map Models.
The principled objective for learning the general flow map ${\bm{\Psi}}_{t\to s}$ is to train a neural network ${\mathbf{G}}_{{\bm{\theta}}}({\mathbf{x}}_{t},t,s)$ by minimizing Equation 4:
@@ Appendix A Related Work
Equations 8 and 9 show that the two parameterizations are tightly connected. In particular,
@@
Hence, the CTM and MF training losses are fundamentally equivalent, differing only by a multiplicative constant. Moreover, in both cases setting $s=0$ recovers the CM scenario, where each state is mapped directly to the clean data. Based on this observation, we will focus our theoretical analysis on ${\bm{\Psi}}_{s\to 0}$ mostly (Appendix F), noting that the same arguments extend naturally to the general ${\bm{\Psi}}_{s\to t}$, including the MF case.
@@ Appendix B Experimental Details
B.1 CIFAR-10, AFHQv2, and FFHQ

### v2

@@ 4.2 Mid-Training with CMT Improves Flow Map Post-Training
These results show that diffusion initialization alone is insufficient and that mid-training is crucial. Using B/4 for pre-training but XL/2 for mid-training further demonstrates that CMT’s mid-training is architecture-agnostic, since it directly learns a map aligned with the teacher sampler.
Table 5: Comparison between MF (initialized with random weights or pretrained SiT) and CMT (MF initialized with CMT mid-trained weights) on the MS-COCO text-to-image (T2I) dataset.
Method Pre-Training Mid-Training Post-Training Total Time ($\downarrow$) 1/2-Step FID ($\downarrow$)
MF (Rand. Init.) 0 0 60 hours 60 hours 15.97 / 5.42
MF (SiT Init.) 22 hours 0 38 hours 60 hours 15.55 / 5.26
CMT 22 hours 1 hours 9 hours 32 hours 15.12 / 5.01
MS-COCO Text-to-Image (T2I): CMT Accelerates and Improves T2I Generation. We follow the setup of U-ViT (Bao et al., 2023) and train MF with the MM-DiT architecture (Esser et al., 2024) on MS-COCO (Lin et al., 2014) from scratch, and we use the validation set for FID evaluation. We consider three MF configurations: (1) MF (Rand. Init.): MF is trained only in the post-training stage, starting from random initialization; (2) MF (SiT Init.): we first pre-train a SiT model and then post-train MF initialized from the SiT weights; (3) CMT: we pre-train SiT, use it as the teacher sampler in the CMT mid-training stage to generate ODE trajectories for an MF model initialized from SiT, and finally post-train MF starting from this mid-trained initializer. Detailed hyperparameters and training settings are provided in Section B.6, and largely follow the ImageNet $256\times 256$ configuration, since MS-COCO shares the same resolution.
Table 5 shows that CMT also performs well for T2I, with significantly faster convergence: it achieves the best FID while reducing total training time by about $47\%$. CMT’s 2-step FID (5.01) is also close to the teacher SiT’s 50-step FID (4.73). We further note that the relatively large 1-step FIDs observed across all variants may be attributable to limitations of the dataset rather than to CMT itself: within a fixed backbone, CMT consistently improves performance and convergence speed over alternative initialization schemes. CMT should therefore be viewed as a general recipe for achieving faster and more stable convergence, not as a modification of the backbone architecture. Together with the ImageNet $256\times 256$ results, these experiments demonstrate that both a smaller MF and a SiT diffusion model can serve as CMT’s teacher and yield strong performance even in T2I generation, highlighting the generality and effectiveness of CMT across problem settings and model architectures.
4.3 Empirical Analysis of the Proposed Mid-Training Scheme
Importance of Mid-Training. We further assess the importance of our mid-training stage on CIFAR-10 using ECT as the flow map model of ${\bm{\Psi}}_{t\to 0}$. To fairly isolate CMT’s mid-training, post-training and vanilla ECT share the same hyperparameters (constant $\Delta t=1/256$, learning rate $10^{-4}$), with all other settings held fixed. We evaluate:
  • Vanilla ECT: Under a 51.2Mimgs budget, we obtain the final 1-step / 2-step FID as 3.54 / 2.12.
  • CMTshort: Short mid-training for 1.28Mimgs + long post-training for 49.92Mimgs. We obtain the final 1-step / 2-step FID as 3.42 / 2.11.
  • CMTlong: Long mid-training for 25.6Mimgs + short post-training for 25.6Mimgs. We obtain the final 1-step / 2-step FID as 3.30 / 2.04.
These results show that longer mid-training outperforms shorter mid-training, which in turn improves over no mid-training, further confirming the importance of CMT ’s mid-training stage.
Ablation Study on Potential Alternatives for Post-Training Initialization. Although we have demonstrated the efficiency of CMT as a mid-training scheme, other variants are possible. We focus on learning the flow map ${\bm{\Psi}}_{t\to 0}$ and highlight two straightforward alternatives. Following our mid-training design principles, these variants aim to be stable and easy to train (e.g., stop-gradient-free regression targets) while using fewer ad-hoc tricks and hyperparameters:
$\mathcal{L}_{\mathrm{var}}^{(1)}({\bm{\theta}}):=\mathbb{E}_{t,\,p_{t}}\!\left[\|{\mathbf{f}}_{\bm{\theta}}({\mathbf{x}}_{t},t)-\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})\|_{2}^{2}\right],\quad\mathcal{L}_{\mathrm{var}}^{(2)}({\bm{\theta}}):=\mathbb{E}_{{\mathbf{x}}_{T}\sim p_{\mathrm{prior}}}\!\left[\|{\mathbf{f}}_{\bm{\theta}}({\mathbf{x}}_{T},T)-\hat{\mathbf{x}}_{0}({\mathbf{x}}_{T})\|_{2}^{2}\right],$
where $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ denotes the estimate at t=0t{=}0 obtained by running the solver with the pre-trained diffusion model starting from the forward-perturbed sample ${\mathbf{x}}_{t}$, and $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{T})$ is obtained similarly by running backward from a prior sample ${\mathbf{x}}_{T}$. We refer to $\mathcal{L}_{\mathrm{var}}^{(1)}({\bm{\theta}})$ as Slow CMT: it follows the same principle as CMT but uses only the endpoints of the ODE trajectory, ignoring intermediate states, so generating the same amount of training data requires substantially more ODE-solver inference. The loss $\mathcal{L}_{\mathrm{var}}^{(2)}({\bm{\theta}})$ is the standard knowledge distillation (KD) objective of Luhman & Luhman (2021).
We compare CMT’s mid-training with KD and Slow CMT on CIFAR-10, training all three under the same settings and then post-training ECT for flow-map learning for 12.8Mimgs with each initialization. Comparing CMT with KD, CMT’s 1-step / 2-step FID (2.74 / 1.97) is significantly better than KD’s (3.54 / 2.19), confirming the benefit of learning intermediate steps. Comparing CMT with Slow CMT, their FIDs are similar (2.74 / 1.97 vs. 2.75 / 1.98), but Slow CMT’s mid-training costs roughly $3\times$ more GPU time because its regression targets cannot reuse intermediate ODE states.
@@ 5 Theoretical Analysis
In mid–training, CMT learns a reliable proxy of the flow map, which then serves as a well–aligned initialization for post–training. To quantify this effect, we analyze the CM flow map ${\bm{\Psi}}_{t\to 0}$ in the post–training stage; the same reasoning applies to general flow maps ${\bm{\Psi}}_{t\to s}$ such as MF.
We compare how well the surrogate CM objective tracks an oracle objective under three initialization schemes: CMT mid–training (${\bm{\theta}}_{\mathrm{CMT}}$), a pre–trained diffusion model (${\bm{\theta}}_{\mathrm{DM}}$), and random initialization (${\bm{\theta}}_{\mathrm{rand}}$). For simplicity, we consider the squared distance $d({\mathbf{x}},{\mathbf{y}})=\|{\mathbf{x}}-{\mathbf{y}}\|_{2}^{2}$ and a uniform weight $w(t)\equiv 1$. Let $\mathcal{L}_{\mathrm{oracle}\text{-}\mathrm{CM}}({\bm{\theta}})$ and $\mathcal{L}_{\mathrm{CM}}({\bm{\theta}})$ denote the oracle and surrogate CM objectives (see Equations 2 and 3). We define the gradient bias $\mathcal{B}({\bm{\theta}}):=\big\|\nabla_{{\bm{\theta}}}\mathcal{L}_{\mathrm{oracle}\text{-}\mathrm{CM}}({\bm{\theta}})-\nabla_{{\bm{\theta}}}\mathcal{L}_{\mathrm{CM}}({\bm{\theta}})\big\|_{2}^{2}$, and evaluate $\mathcal{B}({\bm{\theta}})$ at the initial parameters ${\bm{\theta}}\in\{{\bm{\theta}}_{\mathrm{CMT}},{\bm{\theta}}_{\mathrm{DM}},{\bm{\theta}}_{\mathrm{rand}}\}$. A smaller $\mathcal{B}({\bm{\theta}})$ means that SGD steps on $\mathcal{L}_{\mathrm{CM}}$ closely follow those on the oracle objective. The following informal result summarizes the worst–case bias for each scheme.
Theorem 5.1 (Informal Bias Comparison).
Fix a tolerance $\varepsilon>0$ and a small time step $\Delta t$. Then:
@@
    CMT Initialization. If $\mathcal{L}_{\mathrm{CMT}\text{-}\mathrm{CM}}({\bm{\theta}}_{\mathrm{CMT}})<\varepsilon$, then $\mathcal{B}({\bm{\theta}}_{\mathrm{CMT}})=\mathcal{O}(\varepsilon+\Delta t^{2})$.
@@
    Diffusion Initialization. If $\mathcal{L}_{\mathrm{DM}}({\bm{\theta}}_{\mathrm{DM}})<\varepsilon$, then $\mathcal{B}({\bm{\theta}}_{\mathrm{DM}})=\mathcal{O}\left(\varepsilon+\Delta t^{2}+\mathbb{E}_{t}[\sigma_{t}^{2}/\alpha_{t}^{2}]\right)+\mathbb{E}_{t,{\mathbf{x}}_{t}}\bigl[\|{\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})-\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]\|_{2}^{2}\bigr].$
@@
    Random Initialization. For random weights ${\bm{\theta}}_{\mathrm{rand}}$, $\mathcal{B}({\bm{\theta}}_{\mathrm{rand}})=\mathcal{O}(1)$.
This comparison explains why CMT’s mid–training provides a particularly robust starting point: it already yields a good proxy to the oracle flow map, so the CM gradient is nearly unbiased up to $\mathcal{O}(\varepsilon+\Delta t^{2})$. In contrast, diffusion–based initialization (Geng et al., 2025b) necessarily incurs additional bias from the forward noising process and from the mismatch between the PF–ODE solution ${\bm{\Psi}}_{t\to 0}({\mathbf{x}}_{t})$ and the posterior mean $\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]$, while random initialization can be arbitrarily far from the oracle. We present the complete and rigorous version of this result in Theorem F.1, which also covers initialization from a trained consistency distillation model. In that case, the bias contains an additional, uncontrolled discrepancy term, and in practice its training often requires extra ad–hoc stabilization, further limiting robustness compared to CMT. We refer to Appendix F for the full bias–variance analysis and the resulting excess–risk guarantees.
@@ 6 Conclusion
Statement on the Use of Large Language Models
This work made use of large language models to assist with proofreading and improving the clarity of the writing. All research ideas, theoretical development, experiments, and coding were carried out solely by the authors.
Statement on Reproducibility
We provide all necessary code and models to reproduce CMT’s results at https://github.com/sony/cmt. Comprehensive details of our experimental configurations are further provided in Appendix B to ensure faithful reproducibility.
Statement on Ethics
As with other generative models, CMT can inadvertently produce harmful or inappropriate content (e.g., violent, deepfakes, or derogatory/NSFW materials). These risks are mitigated by enforcing comprehensive safety policies with automated content screening and moderation pipelines that prevent such outputs.
Appendix A Relationship of Flow Map Models
The principled objective for learning the general flow map ${\bm{\Psi}}_{t\to s}$ is to train a neural network ${\mathbf{G}}_{{\bm{\theta}}}({\mathbf{x}}_{t},t,s)$ by minimizing Equation 4 (Lai et al., 2025):
@@ Appendix A Relationship of Flow Map Models
Equations 9 and 10 show that the two parameterizations are tightly connected. In particular,
@@
Hence, the CTM and MF training losses are mathematically related, differing only by a multiplicative constant. Moreover, in both cases setting $s=0$ recovers the CM scenario, where each state is mapped directly to the clean data. Based on this observation, we will focus our theoretical analysis on ${\bm{\Psi}}_{s\to 0}$ mostly (Appendix F), noting that the same arguments extend naturally to the general ${\bm{\Psi}}_{s\to t}$, including the MF case.
@@ Appendix B Experimental Details
B.1 Summary of the CMT Pipeline and Algorithms
Algorithm 1 CMT Pipeline for Fast Flow Map Learning (CM or MF)
Input:
   flow map type $\mathtt{flow\_map\_type}\in\{\texttt{CM},\texttt{MF}\}$, time grid $0=t_{0}<t_{1}<\cdots<t_{M}=T$,
   pre-trained teacher sampler weights $\bm{\phi}$}, student flow map weights $\bm{\theta}$},
   mid-training learning rate $\eta_{\text{mid}}$, post-training learning rate $\eta_{\text{post}}$.
Stage 1: Pre-training (teacher sampler).
if teacher sampler not already available then
  if $\mathtt{flow\_map\_type}=\texttt{CM}$ then
   Train diffusion model ${\mathbf{D}}_{\bm{\phi}}$ with a standard EDM-style objective.
   Construct PF-ODE sampler (e.g. DPM-Solver++) from ${\mathbf{D}}_{\bm{\phi}}$.
  else if $\mathtt{flow\_map\_type}=\texttt{MF}$ then
   Train a flow-matching model ${\mathbf{v}}_{\bm{\phi}}$ or a smaller MF model ${\mathbf{h}}^{\mathrm{teacher}}_{\bm{\phi}}$ in the usual way.
   Construct a deterministic sampler (e.g., a Heun solver with the PF-ODE determined by ${\mathbf{v}}_{\bm{\phi}}$), or an MF multi-step solver based on ${\mathbf{h}}^{\mathrm{teacher}}_{\bm{\phi}}$.
  end if
end if
Freeze $\bm{\phi}$} and treat the resulting sampler as the teacher sampler $\texttt{TeacherSampler}({\bm{\phi}};\cdot)$.
Stage 2: CMT’s Mid-Training.
Initialize $\bm{\theta}$} (optionally ${\bm{\theta}}\leftarrow{\bm{\phi}}$ when architectures match).
repeat
  Sample prior noise ${\mathbf{x}}_{T}\sim p_{\mathrm{prior}}$.
  Generate teacher reference trajectory
      $\{\hat{{\mathbf{x}}}_{t_{i}}\}_{i=0}^{M}\leftarrow\texttt{TeacherSampler}({\bm{\phi}};{\mathbf{x}}_{T})$.
  if $\mathtt{flow\_map\_type}=\texttt{CM}$ then $\triangleright$ Special flow map ${\bm{\Psi}}_{t\to 0}$
   Compute CMT loss as in Equation 7
       $\mathcal{L}_{\mathrm{CMT}\text{-}\mathrm{CM}}({\bm{\theta}})=\sum_{i=0}^{M}d\bigl({\mathbf{f}}_{{\bm{\theta}}}(\hat{{\mathbf{x}}}_{t_{i}},t_{i}),\hat{{\mathbf{x}}}_{t_{0}}\bigr)$.
  else if $\mathtt{flow\_map\_type}=\texttt{MF}$ then $\triangleright$ General flow map ${\bm{\Psi}}_{t\to s}$
   Compute CMT loss as in Equation 8
       $\mathcal{L}_{\mathrm{CMT}\text{-}\mathrm{MF}}({\bm{\theta}})=\sum_{i>j}\bigl\|{\mathbf{h}}_{{\bm{\theta}}}(\hat{{\mathbf{x}}}_{t_{i}},t_{i},t_{j})-\tfrac{\hat{{\mathbf{x}}}_{t_{i}}-\hat{{\mathbf{x}}}_{t_{j}}}{t_{i}-t_{j}}\bigr\|_{2}^{2}$.
  end if
  Update CMT parameters
      ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{\text{mid}}\nabla_{{\bm{\theta}}}\mathcal{L}_{\mathrm{CMT}\text{-}\mathrm{CM}}({\bm{\theta}})$; or ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{\text{mid}}\nabla_{{\bm{\theta}}}\mathcal{L}_{\mathrm{CMT}\text{-}\mathrm{MF}}({\bm{\theta}})$.
until CMT converges (trajectory-aligned initializer $\bm{\theta}$} obtained)
Stage 3: Post-Training of Flow Map Model.
Use the converged CMT mid-trained weights $\bm{\theta}$} as the initialization of the flow map model.
if $\mathtt{flow\_map\_type}=\texttt{CM}$ then $\triangleright$ CM-style post-training for ${\bm{\Psi}}_{t\to 0}$, e.g. ECT/ECD
  repeat
   Sample training batch (data or noise) and times $t$.
   Compute CM post-training loss $\mathcal{L}_{\text{CM}}({\bm{\theta}})$ as in Equation 3.
   Update ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{\text{post}}\nabla_{{\bm{\theta}}}\mathcal{L}_{\text{CM}}({\bm{\theta}})$.
  until post-training converges
else if $\mathtt{flow\_map\_type}=\texttt{MF}$ then $\triangleright$ MF post-training for general ${\bm{\Psi}}_{t\to s}$
  repeat
   Sample training batch and pairs $(t,s)$ with $t>s$.
   Compute MF post-training loss $\mathcal{L}_{\text{MF}}({\bm{\theta}})$ as in Equation 5.
   Update ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{\text{post}}\nabla_{{\bm{\theta}}}\mathcal{L}_{\text{MF}}({\bm{\theta}})$.
  until post-training converges
end if
Output: Learned few-step flow map ${\bm{\Psi}}_{t\to 0}$ (CM) or ${\bm{\Psi}}_{t\to s}$ (MF) with parameters $\bm{\theta}$}.
B.2 CIFAR-10, AFHQv2, and FFHQ